OpenStudy (anonymous):

What is the length of a side of a cube with a volume of 343 cm3

OpenStudy (anonymous):

A cube must have sides of the same length, so the cubed root of the area will give you the length of any side of a cube. \[\sqrt[3]{343}\]

OpenStudy (anonymous):

thank you so much i needed it so bad

OpenStudy (jdoe0001):

 \(\bf volume = x\cdot x\cdot x=x^3\qquad \qquad volume = 343\qquad thus\\ \quad \\ 343=x^3\implies \large \sqrt[3]{343}=x\)